OpenStudy (anonymous):

Lyra and Donna are testing the two-way radios they built for their high school science project. Lyra goes to the top of a building that is 22 meters high with one of the radios and stands directly above the building’s entrance. Donna walks out of the entrance, perpendicular to the building, with the other radio. If Donna is more than 50 meters from the entrance, the connection between the radios breaks. What is the range of communication for the two radios to the nearest meter? 50 meters 55 meters 62 meters 72 meters

OpenStudy (anonymous):

ill give you a medal.

OpenStudy (anonymous):

First, set up a diagram as shown.  You want to assume the maximum (50m) since you are trying to find the range of communication: as in, the MAXIMUM range of communication. Now simply use the pythagorean theorem to find the hypotenuse. Here, "c" represents the hypotenuse. Keep in mind that the hypotenuse represents the range. \[a^2+b^2=c^2\] \[22^2+50^2=c^2\] \[\sqrt{22^2+50^2}=\sqrt{c}\] \[54.6=c\]